) The CPI in 2017 is 252. In 1988, it was 116. What

was the average annual rate of inflation over this 30-

year period?

5.) Suppose when you were born your grandparents

opened a trust fund for you with $30,000 in it. If the

bank pays an average interest rate of 5% annually,

how much money will you have in your trust fund

when you are 21-year-old and have access to it?
Answer #1

ANSWER:

1) Cpi in 2017 = 252

cpi in 1988 = 116

n = 29 yrs (2017 - 1988)

inflation = ?

inflation = (cpi in 2017 / cpi in 1988) ^ 1/n - 1

inflation = (252 / 116) ^ 1 / 29 - 1

inflation = (2.1724) ^ 1 / 29 - 1

inflation = 1.027114 - 1

inflation = 0.027114 or 2.71%

2) present value = $30,000

i = 5%

n = 21

fv = ?

fv = pv(f/p,i,n)

fv = 30,000(f/p,5%,21)

fv = 30,000 * 2.786

fv = 83,580

so the value of the trust fundd after 21 yrs is $83,580.